Mark works strictly on commission from his gross sales selling 2 different products for his company. Last month, his gross sales were $88,000. If he earns 6% commission on product A, and 4% commission on product B, what were his gross sales for each product if he earned $4880 in total commission?
Answer by richwmiller(17219)

Total amount of sales: $88000
x+y=88000,
Total commission for the two products is: $4880
0.06*x+0.04*y=4880
x=88000-y
Substitute for x
0.06*(88000-y)+0.04*y=4880
Multiply out
5280-0.06*y+0.04*y=4880
Combine like terms.
-0.02*y=-400
Isolate y
y=$ 20000.00 at 4%
x=88000-y
Calculate x
x=$ 68000.00 at 6%
Check
0.06*68000+0.04*20000=4880
4080+800=4880
4880=4880
If this statement is TRUE and neither x nor y is negative then all is well
code int1